OpenStudy (anonymous):

Choose the correct slope of the line that passes through the points (1, −3) and (3, −5).

OpenStudy (anonymous):

here are the possible answers 1) x-intercept: (0, −3), y-intercept: (0, 2) 2) x-intercept: (−3, 0), y-intercept: (2, 0) 3) x-intercept: (0, 2), y-intercept: (0, −3) 4) x-intercept: (2, 0), y-intercept: (0, −3)

OpenStudy (aravindg):

The qn asks for slope and the options are intercepts ? o.O

OpenStudy (anonymous):

equation is y = -x -2 ... slope would be -1.. y int is -2. I don't see it fitting into your options though

OpenStudy (anonymous):

oh sorry wait

OpenStudy (anonymous):

−1 1 0 1 over 2

OpenStudy (anonymous):

The answer is the first choice then :)

OpenStudy (aravindg):

slope=y difference/x difference

OpenStudy (aravindg):

for points (1, −3) and (3, −5). y diff=-5+3 z diff=3-1

OpenStudy (aravindg):

slope =-2/2=-1